# Quasiparticle interference and spectral function of the UTe2 superconductive surface band

### v1

@@ Abstract
We compute the (0-11) surface spectral function, the surface density of states (DOS), and the quasiparticle interference (QPI) patterns, both in the normal state and superconducting (SC) state of UTe2. We consider all possible non-chiral and chiral order parameters (OPs) that could in principle describe the superconductivity in this compound. We describe the formation of surface states whose maximum intensity energy depends on the nature of the pairing. We study also the QPI patterns resulting from the scattering of these surface states. Along the lines of Ref. 1, we show that the main feature distinguishing between various OPs is a QPI peak that is only observed experimentally in the superconducting state. The energy dispersion and the stability of this peak is consistent among the non-chiral OPs only with a $B_{3u}$ pairing. Moreover, $B_{3u}$ is the only non-chiral pairing that shows a peak at zero energy in the DOS, consistent with the experimental observations.
@@ I Introduction
The heavy-fermion material uranium ditelluride (UTe2) has recently been identified as a superconductor [2, 3, 4, 5, 6], with a critical temperature of $T_{c}\sim 1.6-2.0$ K. Notably, unlike other uranium-based compounds [7, 8], superconductivity in UTe2 emerges from a paramagnetic normal state rather than from an ordered magnetic phase. Beyond its superconducting properties, UTe2 displays other intriguing phenomena, including charge density wave (CDW) order [9, 10], which coexists with superconductivity—though this coexistence may be restricted to the (0-11) surface [11, 12]. This interaction may lead to the formation of a modulated superconducting state [13], suggesting the possible emergence of a pair-density wave (PDW) state [9].
@@
Advancing our understanding of the underlying physics of UTe2 relies on determining the symmetry of its superconducting order parameter (OP). While this is a well-defined and fundamental question, it has proven extremely challenging to resolve in non-BCS superconductors, as evidenced by the long-standing controversies surrounding Sr2RuO4 [14, 15]. UTe2 is widely believed to be a triplet superconductor, supported by various experimental observations, including an upper critical field $H_{c2}$ that significantly exceeds the Pauli limit [4, 3] and a minimal change in the Knight shift upon entering the superconducting phase [5, 3], among other indications. The literature seems to converge on a $p$-wave-type pairing, with no argument for an $f$-wave pairing, consistent also with the fact that higher angular momentum pairings are in general energetically disfavored, .
@@
However, there is far less consensus regarding the specific odd-parity representation of the superconducting order parameter. Soon after the discovery of the superconducting phase, observations of two distinct superconducting transitions in the specific heat [16], power law dependence of the magnetic penetration depth [17], along with a nonzero polar Kerr effect [16] indicating time-reversal symmetry breaking (T), suggested that UTe2 could be a two-component chiral superconductors [18] with chiral edge states [19]. However more recent measurements with better quality sample now seem to point to a single component superconductor. This is supported by recent reports showing no evidence of broken time-reversal symmetry, as indicated by the absence of a polar Kerr effect [20], along with zero-energy Andreev peak measurements [21]. Additional confirmation comes from ultrasound spectroscopy [22] and SQUID measurements [23], reinforcing the case for a single-component superconducting order parameter.
@@
For a single-component order parameter, there are four possible pairing symmetries, corresponding to the odd-parity irreducible representations of the point group $D_{2h}$ of UTe2222, namely ${A_{u},B_{1u},B_{2u},B_{3u}}$, with $A_{u}$ and $B_{1u}$ corresponding to gapped order parameters, while $B_{2u}$ and $B_{3u}$ are gapless. There is no consensus on which pairing symmetry is realized in UTe2222, with studies suggesting all possibilities, $A_{u}$ [24, 25], $B_{1u}$[23], $B_{2u}$[22], and $B_{3u}$ [26, 23]. The ongoing debate surrounding the determination of the superconducting order parameter symmetry in UTe2, driven by various experimental observations, highlights the necessity of employing diverse experimental and theoretical approaches to make meaningful progress.
@@
One popular approach that helps give insight into the nature of the superconducting order parameter is the study of the modifications to the density of states near an impurity. This can be directly probed using scanning tunneling microscopy (STM). The Fourier transforms of the real space STM maps in the presence of impurities are also denoted quasiparticle interference (QPI) patterns. This method has been successfully applied to various superconductors, including $d$-wave cuprates [27], iron-pnictides [28, 29], and strontium ruthenates [30]. However, its main limitation, especially for bulk 3D materials, is that it serves only as a surface probe.
@@
In this work, along the lines of Ref. 1, we investigate how distinct surface QPI patterns arising from different pairing symmetries can serve as a diagnostic tool to distinguish between various superconducting states in UTe2. We use a recently developed Green’s function technique [31, 32] to compute the surface Green’s function for the experimentally relevant (0-11) surface, as UTe2 cleaves easily along this plane. Subsequently, using the well-established T-matrix formalism [33], we compute the QPI patterns generated by a surface-localized impurity using the (0-11) surface Green’s function.
@@
Several minimal models [34, 35, 36] have been proposed based on ab-initio computations to explain the normal-state properties of UTe2. It is widely believed that the Fermi surface of UTe2 consists of two cylindrical sheets extending along the $c$-axis [37], with only weak dispersion in this direction as measured by magnetic quantum oscillations [38]. Some authors have also proposed the existence of a small 3D electron pocket, originating from strongly correlated $f$-electrons [39, 40], centered around the $\Gamma$-point. However, since no definitive experimental signature has been observed to date, we will not explore this possibility in the current work. Regarding the superconducting properties, although a single-component order parameter appears more likely, we focus on studying both single-component (${A_{u},B_{1u},B_{2u},B_{3u}}$) and multi-component triplet pairing symmetries ($A_{u}+iB_{1u}$, $A_{u}+iB_{2u}$, $A_{u}+iB_{3u}$, $B_{1u}+iB_{2u}$, $B_{1u}+iB_{3u}$, $B_{2u}+iB_{3u}$).
@@
While the main conclusions of our study have already been presented in Ref. 1, in the present work we provide a detailed development of the theoretical framework, as well as a comprehensive description of the theoretical analysis supporting the experimental STM results previously reported. Also we extend the analysis of Ref. 1, which was focused mainly on the non-chiral $B_{2u}$ and $B_{3u}$ pairings to all the possible chiral and non-chiral pairing symmetries.
@@
Our results indicate the formation of surface states for all the order parameters studied, however the only one showing a maximum of intensity at zero energy is $B_{3u}$, the rest appearing to be centered rather at finite energies, whose values depend on the form of the OP. Consequently the surface DOS will have a peak at zero energy only for $B_{3u}$ and will show a double peak with a dip in the middle for all the rest of the OPs, however, due to the large quasiparticle damping, the split of the central peak will not visible for all the OPs. All the surface states seem to have a maximum of intensity in the same regions in momentum space, following the normal state surface state pattern. However, at a given energy the resulting QPI differ greatly from one order parameter to the next. This is due first to the intensity distribution with energy which is different for each OP, but most importantly to the destructing interference stemming from phase cancellations which makes certain features disappear completely for the normal state or some of the OPs. We show that the spin structure of the surface states may play a role in these phase cancellations, however it is general hard to predict which wavevectors will be attenuated in the QPI patterns.
@@
The structure of the paper is the following. In Section II, we outline the tight-binding model and the method used to compute the surface spectral function and the QPI patterns for an impurity localized on the (0-11) surface of UTe2. In Sections III and IV, we present the results obtained for UTe2 in its normal state, and respectively in its superconducting state, for both non-chiral and chiral pairings. We discuss these results and conclude in Section V. Additional details on the model and method are provided in Appendix A, and a detailed discussion on the symmetry constraints is given in Appendix B. In Appendices D and C we provide the plots for the bulk spectral function and the JDOS at zero-energy, as well as the plots for the surface spectral function and the QPI patterns at finite energies.
@@ II.1 Tight-binding model
Figure 1: Schematic picture of the UTe2 lattice. The 4-orbital tight-binding model used in this work includes nearest-neighbor hopping between U atoms, nearest-neighbor hopping between blue color Te atoms, as gray color Te atoms weakly affect the Fermi surface[22], and hybridization $\delta$ between U and Te energy bands.
UTe2 has a body-centered orthorhombic lattice structure with the space group symmetry Immm [41] and lattice constants $a=0.41$ nm, $b=0.61$ nm and $c=1.39$ nm for the dimensions of the unit cell along $\widehat{x}$-, $\widehat{y}$- and $\widehat{z}$-axis (see Fig. 1). We use a 4-orbital tight-binding model based on density functional theory (DFT) calculations introduced by Theuss et al. [22]. Two sets of parameters are given by these authors: one matching DFT results and the other matching quantum-oscillation (QO) experiments. Our approach was to start with the DFT tight-binding parameters and slightly modify their values (see Appendix A.1) to match the experimental constraints on the Fermi surface. This model, which exhibits two cylindrical-like Fermi surface sheets, does not consider the hypothesized Fermi pocket centered around $\Gamma$ predicted by some authors based on strongly correlated physics beyond DFT-based approaches [39], due to conflicting experimental evidence [40, 38]. The presence of two uranium atoms and two tellurium atoms per unit cell leads to four bands of energy described by the Hamiltonian written in the orbital space $\mathcal{E}_{orb}$
$\mathcal{H}_{\mathrm{TB}}(\mathbf{k})=\begin{pmatrix}\mathcal{H}_{\mathrm{U}}(\mathbf{k})&\mathcal{H}_{\delta}\\ \mathcal{H}_{\delta}^{\dagger}&\mathcal{H}_{\mathrm{Te}}(\mathbf{k})\\ \end{pmatrix}$ (1)
@@
$\mathcal{H}_{\mathrm{U}}(\mathbf{k})$ $=$ $\begin{pmatrix}\mathcal{E}_{\mathrm{U}}(\mathbf{k})&f_{\mathrm{U}}(\mathbf{k})\\ f_{\mathrm{U}}^{*}(\mathbf{k})&\mathcal{E}_{\mathrm{U}}(\mathbf{k})\\ \end{pmatrix}$ (2)
$\mathcal{H}_{\mathrm{Te}}(\mathbf{k})$ $=$ $\begin{pmatrix}\mathcal{E}_{\mathrm{Te}}(\mathbf{k})&f_{\mathrm{Te}}(\mathbf{k})\\ f_{\mathrm{Te}}^{*}(\mathbf{k})&\mathcal{E}_{\mathrm{Te}}(\mathbf{k})\\ \end{pmatrix}\;$ (3)
@@
$\mathcal{E}_{\mathrm{U}}(\mathbf{k})$ $=$ $\mu_{U}-2t_{U}\cos(ak_{x})-2t_{ch,U}\cos(bk_{y})$ (4)
@@
-- $4t_{z,U}e^{-ick_{z}/2}\cos(ak_{x}/2)\cos(bk_{y}/2)$
$\mathcal{E}_{\mathrm{Te}}(\mathbf{k})$ $=$ $\mu_{T_{e}}-2t_{ch,T_{e}}\cos(ak_{x})$ (6)
@@
$\mathcal{H}_{\delta}=\begin{pmatrix}\delta&0\\ 0&\delta\\ \end{pmatrix}$ (8)
[…]
@@ II.2 Expression for the gap function
In this work we restrict the analysis to a triplet $p$-wave superconducting whose possible physical origin is the pairing driven by a ferromagnetic quantum critical point [3, 19]. Pairing symmetries in a crystal are classified by the point group of the model: for UTe2 the point group is $D_{2h}$ [36]. Because inversion symmetry $\mathcal{I}\in D_{2h}$, the triplet pairing can be classified either as odd or as even under inversion. The gap function for a triplet superconductor in spin space can be written as:
$\Delta(\mathbf{k})=\left[{\bf d}(\mathbf{k})\cdot\boldsymbol{\sigma}\right]\left(i\sigma_{y}\right),$ (9)
@@
with $\boldsymbol{\sigma}=\left\{\sigma_{x},\sigma_{y},\sigma_{z}\right\}$ the Pauli matrix, and ${\bf d}(\mathbf{k})$ a vector which characterizes the pairing. In single-orbital superconductors, the Pauli principle implies that the gap function should satisfy the relation $\Delta(\mathbf{k})=-\Delta^{T}(-\mathbf{k})$. For more complicated orbital pairings, such as the ones we study here, $\Delta(\mathbf{k})$ can also be non-trivial in orbital space. Because the electrons contributing to the Fermi surface are $f$-electrons, they experience a strong spin-orbit coupling. Consequently, point group transformations and spin transformations are no longer independent: acting with certain symmetries on the lattice results in corresponding actions on the spin [43]. This implies that the direction of the ${\bf d}(\mathbf{k})$ vector is constrained by the point group representation $D_{2h}$ which has eight irreducible representations: $A_{g/u}$, $B_{1g/u}$, $B_{2g/u}$ and $B_{3g/u}$ [44]. However, since we consider triplet pairing, it should be odd under inversion $\mathcal{I}$, such that the allowed representations are $A_{u}$, $B_{1u}$, $B_{2u}$ and $B_{3u}$. All these representations correspond to non-chiral order parameters invariant under the time-reversal symmetry represented by $\mathcal{T}=i\sigma_{y}\mathcal{K}$, where $\mathcal{K}$ is complex conjugation. Chiral order parameters [45] can be constructed by adding two representations which are energetically degenerate with a relative phase of $\pi/2$ such that it breaks time-reversal symmetry [46]. Since we are not solving the self-consistent BCS gap equations to compare the free energies, we consider different combinations even if they may be disfavored in reality. The transformation properties of these pairing symmetries are listed in Tab. 1.
@@
Because there are two uranium orbitals and two tellurium orbitals in our tight-binding model, we can have both intra- and inter-orbital pairing. In the following we only consider inter-orbital pairing states with orbital-triplet pairings [36] and restrict the pairings to nearest-neighbour couplings along the $\widehat{x}$-, $\widehat{y}$- or $\widehat{z}$-axis. We moreover assume that the pairing strength is independent of the axis direction. This lead to the following form for the non-chiral ${\bf d}$ vector
${\bf d}_{A_{u}}(\mathbf{k})$ $=$ $\Delta_{0}\,\Big{(}\sin(ak_{x}),\sin(bk_{y}),\sin(ck_{z})\Big{)}^{T}$ (10)
${\bf d}_{B_{1u}}(\mathbf{k})$ $=$ $\Delta_{0}\,\Big{(}\sin(bk_{y}),\sin(ak_{x}),0\Big{)}^{T}$ (11)
${\bf d}_{B_{2u}}(\mathbf{k})$ $=$ $\Delta_{0}\,\Big{(}\sin(ck_{z}),0,\sin(ak_{x})\Big{)}^{T}$ (12)
${\bf d}_{B_{3u}}(\mathbf{k})$ $=$ $\Delta_{0}\,\Big{(}0,\sin(ck_{z}),\sin(bk_{y})\Big{)}^{T}$ (13)
@@
where $\Delta_{0}$ is the pairing strength. The form of the chiral ${\bf d}$ vectors follow readily by adding two non-chiral ${\bf d}$ vectors with a relative phase of $\pi/2$. Note that the point group $D_{2h}$ does not impose that pairing strength $\Delta_{0}$ has to be independent of the axis direction, meaning that one can relax this assumption and take a more general ${\bf d}(\mathbf{k})$ if needed.
@@
$B_{2u}+iB_{3u}$1 $\times$ $\times$ 1 $\times$ $\times$

### v2

@@ Abstract
We compute the (0-11) surface spectral function, the surface density of states (DOS), and the quasiparticle interference (QPI) patterns, both in the normal state and superconducting (SC) state of UTe2. We consider all possible non-chiral and chiral order parameters (OPs) that could in principle describe the superconductivity in this compound. We describe the formation of surface states whose maximum intensity energy depends on the nature of the pairing. We also study the QPI patterns resulting from the scattering of these surface states. Along the lines of [Nat. Phys. 21, 1555 (2025)], we show that the main feature distinguishing between various OPs is a QPI peak that is only observed experimentally in the superconducting state. The energy dispersion and the stability of this peak is consistent among the non-chiral OPs only with a $B_{3u}$ pairing. Moreover, $B_{3u}$ is the only non-chiral pairing that shows a peak at zero energy in the DOS, consistent with the experimental observations.
@@ I Introduction
The heavy-fermion material uranium ditelluride (UTe2) has recently been identified as a superconductor [aoki2019unconventional, ran2019nearly, knebel2019field, nakamine2019superconducting, tokunaga2019125te], with a critical temperature of $T_{c}\sim 1.6-2.0$ K. Notably, unlike other uranium-based compounds [saxena2000superconductivity, aoki2001coexistence], superconductivity in UTe2 emerges from a paramagnetic normal state rather than from an ordered magnetic phase. Beyond its superconducting properties, UTe2 displays other intriguing phenomena, including charge density wave (CDW) order [aishwarya2023magnetic, lafleur2024inhomogeneous], which coexists with superconductivity—though this coexistence may be restricted to the (0-11) surface [theuss2024absence, kengle2024absence]. This interaction may lead to the formation of a modulated superconducting state [gu2023detection], suggesting the possible emergence of a pair-density wave (PDW) state [aishwarya2023magnetic].
@@
Advancing our understanding of the underlying physics of UTe2 relies on determining the symmetry of its superconducting order parameter (OP). While this is a well-defined and fundamental question, it has proven extremely challenging to resolve in non-BCS superconductors, as evidenced by the long-standing controversies surrounding Sr2RuO4 [mackenzie2017even, pustogow2019constraints]. UTe2 is widely believed to be a triplet superconductor, supported by various experimental observations, including an upper critical field $H_{c2}$ that significantly exceeds the Pauli limit [knebel2019field, ran2019nearly] and a minimal change in the Knight shift upon entering the superconducting phase [nakamine2019superconducting, ran2019nearly], among other indications. However, more recent experiments suggest that these results are dependent on the crystallographic axis along which the magnetic field is applied [matsumura2023large]. The literature seems to converge on a $p$-wave-type pairing, with no argument for an $f$-wave pairing, consistent also with the fact that higher angular momentum pairings are in general energetically disfavored.
@@
However, there is far less consensus regarding the specific odd-parity representation of the superconducting order parameter. Soon after the discovery of the superconducting phase, observations of two distinct superconducting transitions in the specific heat [hayes2021multicomponent], power law dependence of the magnetic penetration depth [ishihara2023chiral], along with a nonzero polar Kerr effect [hayes2021multicomponent] indicating time-reversal symmetry breaking (T), suggested that UTe2 could be a two-component chiral superconductors [kittaka2020orientation] with chiral edge states [jiao2020chiral]. However, more recent measurements with better quality sample now seem to point to a single component superconductor. This is supported by recent reports showing no evidence of broken time-reversal symmetry, as indicated by the absence of a polar Kerr effect [ajeesh2023fate], along with zero-energy Andreev peak measurements [gu2025pair]. Additional confirmation comes from ultrasound spectroscopy [theuss2024single] and SQUID measurements [iguchi2023microscopic], reinforcing the case for a single-component superconducting order parameter.
@@
For a single-component order parameter, there are four possible pairing symmetries, corresponding to the odd-parity irreducible representations of the point group $D_{2h}$ of UTe2, namely $A_{u},B_{1u},B_{2u},B_{3u}$}. The gapped or gapless nature of these superconducting order parameters depends on the topology of the normal-state Fermi surface of UTe2, which remains a subject of active debate. Within the model considered in this work, the $A_{u}$ and $B_{1u}$ order parameters are fully gapped, whereas the $B_{2u}$ and $B_{3u}$ states exhibit gapless behavior. There is no consensus on which pairing symmetry is realized in UTe2, with studies suggesting all possibilities, $A_{u}$ [matsumura2023large, suetsugu2024fully], $B_{1u}$[iguchi2023microscopic], $B_{2u}$[theuss2024single], and $B_{3u}$ [fujibayashi2022superconducting, iguchi2023microscopic]. The ongoing debate surrounding the determination of the superconducting order parameter symmetry in UTe2, driven by various experimental observations, highlights the necessity of employing diverse experimental and theoretical approaches to make meaningful progress.
@@
One popular approach that helps give insight into the nature of the superconducting order parameter is the study of the modifications to the density of states near an impurity. This can be directly probed using scanning tunneling microscopy (STM). The Fourier transforms of the real-space STM maps in the presence of impurities are also denoted quasiparticle interference (QPI) patterns. This method has been successfully applied to various superconductors, including $d$-wave cuprates [hudson2001interplay], iron-pnictides [allan2012anisotropic, sprau2017discovery], and strontium ruthenates [sharma2020momentum]. However, its main limitation, especially for bulk 3D materials, is that it serves only as a surface probe.
@@
In this work, along the lines of Ref. experimentalpaper, we investigate how distinct surface QPI patterns arising from different pairing symmetries can serve as a diagnostic tool to distinguish between various superconducting states in UTe2. We use a recently developed Green’s function technique [kaladzhyan2019obtaining, pinon2020surface] to compute the surface Green’s function for the experimentally relevant (0-11) surface, as UTe2 cleaves easily along this plane. Subsequently, using the well-established T-matrix formalism [balatsky2006impurity], we compute the QPI patterns generated by a surface-localized impurity using the (0-11) surface Green’s function.
@@
Several minimal models [ishizuka2021periodic, shishidou2021topological, tei2023possible] have been proposed based on ab-initio computations to explain the normal-state properties of UTe2. It is widely believed that the Fermi surface of UTe2 consists of two cylindrical sheets extending along the $c$-axis [xu2019quasi], with only a weak dispersion in this direction as measured by magnetic quantum oscillations [eaton2024quasi]. Some authors have also proposed the presence of a small three-dimensional electron pocket originating from strongly correlated $f$-electrons, located either near the $\Gamma$-point [choi2022correlated, broyles2023revealing] or the $Z$-point [miao2020low]. This hypothesis is supported among others by core-level spectroscopy indirect evidence from transport measurements [Fujimori2019, fujimori2021core, eo2022c]. Although the presence of these Fermi pockets would influence the QPI interference pattern, since no definitive experimental signature has been observed to date, we will not explore this possibility in the current work. Regarding the superconducting properties, although a single-component order parameter appears more likely, we focus on studying both single-component ($A_{u},B_{1u},B_{2u},B_{3u}$}) and multi-component ($A_{u}+iB_{1u}$, $A_{u}+iB_{2u}$, $A_{u}+iB_{3u}$, $B_{1u}+iB_{2u}$, $B_{1u}+iB_{3u}$, $B_{2u}+iB_{3u}$) triplet pairing symmetries.
@@
Although the main conclusions of our study have already been presented in Ref. experimentalpaper, in the present work we provide a detailed development of the theoretical framework, as well as a comprehensive description of the theoretical analysis supporting the experimental STM results previously reported. Also, we extend the analysis of Ref. experimentalpaper, which was focused mainly on the non-chiral $B_{2u}$ and $B_{3u}$ pairings to all the possible chiral and non-chiral pairing symmetries.
@@
Our results indicate the formation of surface states for all the order parameters studied, however the only one showing a maximum of intensity at zero energy is $B_{3u}$, the rest appearing to be centered rather at finite energies, whose values depend on the form of the OP. Consequently, the surface DOS will have a peak at zero energy only for $B_{3u}$ and will show a double peak with a dip in the middle for all the rest of the OPs, however, due to the large quasiparticle damping, the split of the central peak will not be visible for all the OPs. All the surface states seem to have a maximum of intensity in the same regions in momentum space, following the normal state surface state pattern. However, at a given energy the resulting QPI differ greatly from one order parameter to the next. This is due first to the intensity distribution with energy which is different for each OP, but most importantly to the destructing interference stemming from phase cancellations which makes certain features disappear completely for the normal state or some of the OPs. We show that the spin structure of the surface states may play a role in these phase cancellations, however it is general hard to predict which wavevectors will be attenuated in the QPI patterns.
@@
The structure of the paper is the following. In Section II, we outline the tight-binding model and the method used to compute the surface spectral function and the QPI patterns for an impurity located on the (0-11) surface of UTe2. In Sections III and IV, we present the results obtained for UTe2 in its normal state, and respectively in its superconducting state, for both non-chiral and chiral pairings. We discuss these results and conclude in Section V. Additional details on the model and method are provided in Appendix A, and a detailed discussion on symmetry constraints is given in Appendix B. In Appendices  C and D we provide the plots for the bulk spectral function and the JDOS at zero-energy, as well as the plots for the surface spectral function and the QPI patterns at finite energies.
@@ II.1 Tight-binding model
Figure 1: Schematic picture of the UTe2 lattice. The 4-orbital tight-binding model used in this work includes nearest-neighbor hopping between U atoms, nearest-neighbor hopping between blue color Te atoms, as gray color Te atoms weakly affect the Fermi surface[theuss2024single], and hybridization $\delta$ between U and Te energy bands.
UTe2 has a body-centered orthorhombic lattice structure with the space group symmetry Immm [bradley2009mathematical] and lattice constants $a=0.41$ nm, $b=0.61$ nm and $c=1.39$ nm for the dimensions of the unit cell along $\widehat{x}-$, $\widehat{y}-$ and $\widehat{z}-$axis (see Fig. 1). We use a four-orbital tight-binding model based on density functional theory (DFT) calculations introduced by Theuss et al. [theuss2024single]. Two sets of parameters are given by these authors: one matching DFT results and the other matching quantum-oscillation (QO) experiments. Our approach was to start with the DFT tight-binding parameters and modify their values (see Appendix A.1) to match the experimental constraints on the Fermi surface. This model, which exhibits two cylindrical-like Fermi surface sheets, does not consider the hypothesized Fermi pocket centered around $\Gamma$ predicted by some authors based on strongly correlated physics beyond DFT-based approaches [choi2022correlated], due to conflicting experimental evidence [broyles2023revealing, eaton2024quasi]. The presence of two uranium atoms and two tellurium atoms per unit cell leads to four bands of energy described by the Hamiltonian written in orbital space $\mathcal{E}_{orb}$
$\mathcal{H}_{\mathrm{TB}}(\mathbf{k})=\begin{pmatrix}\mathcal{H}_{\mathrm{U}}(\mathbf{k})&\mathcal{H}_{\delta}\\ \mathcal{H}_{\delta}^{\dagger}&\mathcal{H}_{\mathrm{Te}}(\mathbf{k})\\ \end{pmatrix}$ $~,$ (1)
@@
$\mathcal{H}_{\mathrm{U}}(\mathbf{k})$ $=$ $\begin{pmatrix}\mathcal{E}_{\mathrm{U}}(\mathbf{k})&f_{\mathrm{U}}(\mathbf{k})\\ f_{\mathrm{U}}^{*}(\mathbf{k})&\mathcal{E}_{\mathrm{U}}(\mathbf{k})\\ \end{pmatrix}~$, (2)
$\mathcal{H}_{\mathrm{Te}}(\mathbf{k})$ $=$ $\begin{pmatrix}\mathcal{E}_{\mathrm{Te}}(\mathbf{k})&f_{\mathrm{Te}}(\mathbf{k})\\ f_{\mathrm{Te}}^{*}(\mathbf{k})&\mathcal{E}_{\mathrm{Te}}(\mathbf{k})\\ \end{pmatrix}~$, (3)
@@
$\mathcal{E}_{\mathrm{U}}(\mathbf{k})$ $=$ $\mu_{U}-2t_{U}\cos(ak_{x})-2t_{ch,U}\cos(bk_{y})~,$ (4)
@@
- $4t_{z,U}e^{-ick_{z}/2}\cos(ak_{x}/2)\cos(bk_{y}/2)~,$
$\mathcal{E}_{\mathrm{Te}}(\mathbf{k})$ $=$ $\mu_{T_{e}}-2t_{ch,T_{e}}\cos(ak_{x})~,$ (6)
@@
$\mathcal{H}_{\delta}=\begin{pmatrix}\delta&0\\ 0&\delta\\ \end{pmatrix}~$, (8)
with $\delta$ the hybridization strength between $U$ and $Te$ atoms. In the absence of hybridization, i.e. for $\delta=0$, two of the four bands cross the Fermi energy. Introducing a hybridization creates electron and hole pockets with cylindrical-like shapes aligned along the $\widehat{z}-$axis. We verified that the specific form of the hybridization is not crucial as long as it results in two disconnected pockets. We chose the value of the hybridization parameter $\delta$ such that the area of the electron and hole pockets at $k_{z}=0$ matches recent quantum-oscillation experiments [eaton2024quasi, weinberger2024quantum], with $\mathcal{A}_{e}\approx\mathcal{A}_{h}\approx\ 33.6$ nm2. The resulting Fermi surface is shown in Fig. 2. The main characteristic is the presence of weakly dispersing cylindrical-like Fermi pockets parallel to the $\widehat{z}-$axis . We see moreover that this Fermi surface is symmetrical under the Immm space group symmetries.
@@ II.2 Expression for the gap function
In this work we restrict the analysis to a triplet $p$-wave superconducting whose possible physical origin is the pairing driven by a ferromagnetic quantum critical point [ran2019nearly, jiao2020chiral]. Pairing symmetries in a crystal are classified by the point group of the model: for UTe2 the point group is $D_{2h}$ [tei2023possible]. Because inversion symmetry $\mathcal{I}\in D_{2h}$, the triplet pairing can be classified either as odd or as even under inversion. The gap function for a triplet superconductor in spin space can be written as:
$\Delta(\mathbf{k})=\left[{\bf d}(\mathbf{k})\cdot\boldsymbol{\sigma}\right]\left(i\sigma_{y}\right)~,$ (9)
@@
with $\boldsymbol{\sigma}=\left\{\sigma_{x},\sigma_{y},\sigma_{z}\right\}$ the Pauli matrix, and ${\bf d}(\mathbf{k})$ a vector which characterizes the pairing. In single-orbital superconductors, the Pauli principle implies that the gap function should satisfy the relation $\Delta(\mathbf{k})=-\Delta^{T}(-\mathbf{k})$. For more complicated orbital pairings, such as those we study here, $\Delta(\mathbf{k})$ can also be non-trivial in orbital space. Because the electrons contributing to the Fermi surface are $f$-electrons, they experience a strong spin-orbit coupling. Consequently, point group transformations and spin transformations are no longer independent: acting with certain symmetries on the lattice results in corresponding actions on the spin [sigrist1991phenomenological]. This implies that the direction of the ${\bf d}(\mathbf{k})$ vector is constrained by the point group representation $D_{2h}$ which has eight irreducible representations: $A_{g/u}$, $B_{1g/u}$, $B_{2g/u}$ and $B_{3g/u}$ [aroyo2006bilbao]. However, since we consider triplet pairing, it should be odd under inversion $\mathcal{I}$, such that the allowed representations are $A_{u}$, $B_{1u}$, $B_{2u}$ and $B_{3u}$. All these representations correspond to non-chiral order parameters invariant under the time-reversal symmetry represented by $\mathcal{T}=i\sigma_{y}\mathcal{K}$, where $\mathcal{K}$ is complex conjugation. Chiral order parameters [kallin2016chiral] can be constructed by adding two representations which are energetically degenerate with a relative phase of $\pi/2$ such that it breaks the time-reversal symmetry [sato2017topological]. Since we are not solving the self-consistent BCS gap equations to compare the free energies, we consider different combinations, even if they may be disfavored in reality. The transformation properties of these pairing symmetries are listed in Tab. 1.
@@
Because there are two uranium orbitals and two tellurium orbitals in our tight-binding model, we can have both intra- and inter-orbital pairing. In the following, we only consider inter-orbital pairing states with orbital-triplet pairings [tei2023possible] and restrict the pairings to nearest-neighbor couplings along the $\widehat{x}-$, $\widehat{y}-$ or $\widehat{z}-$axis. We, moreover, assume that the pairing strength is independent of the axis direction. This leads to the following form for the non-chiral d{d}-vector
${\bf d}_{A_{u}}(\mathbf{k})$ $=$ $\Delta_{0}\,\Big(\sin(ak_{x}),\sin(bk_{y}),\sin(ck_{z})\Big)^{T}~,$ (10)
${\bf d}_{B_{1u}}(\mathbf{k})$ $=$ $\Delta_{0}\,\Big(\sin(bk_{y}),\sin(ak_{x}),0\Big)^{T}~,$ (11)
${\bf d}_{B_{2u}}(\mathbf{k})$ $=$ $\Delta_{0}\,\Big(\sin(ck_{z}),0,\sin(ak_{x})\Big)^{T}~,$ (12)
${\bf d}_{B_{3u}}(\mathbf{k})$ $=$ $\Delta_{0}\,\Big(0,\sin(ck_{z}),\sin(bk_{y})\Big)^{T}~,$ (13)
@@
where $\Delta_{0}$ is the pairing strength. The form of the chiral d{d}-vectors follows readily by adding two non-chiral d{d}-vectors with a relative phase of $\pi/2$. Note that the point group $D_{2h}$ does not impose that pairing strength $\Delta_{0}$ has to be independent of the axis direction, meaning that one can relax this assumption and take a more general $\mathbf{d}(\mathbf{k})$ if needed. Others terms are also allowed by symmetry for the d{d}-vectors [hillier2012nonunitary]. For instance, the $d$-vector for $B_{3u}$ pairing can be generalized to $\mathbf{d}_{B_{3u}}(\mathbf{k})=\left(C_{0}\sin(ak_{x})\sin(bk_{y})\sin(ck_{z}),\Delta_{0}\sin(ck_{z}),\Delta_{0}\sin(bk_{y})\right)^{T}$. The QPI results for $C_{0}\leq\Delta_{0}$ are qualitatively unchanged.
@@
$B_{2u}+iB_{3u}$ -1 $\times$ $\times$ 1 $\times$ $\times$